A system at equilibrium is represented by the following chemical equation: N₂(g) + 3H₂(g) <--> 2NH₃(g) At a given temperature, 1.20 moles of N₂, 0.866 moles of H₂, and 4.17 * 10⁻⁸ moles of NH₃ are in the 4.00 L vessel that contains the equilibrium. Calculate the value of K for the system.
1
Expert's answer
2020-06-03T15:17:06-0400

Solution.

"n(N_2)=1.20mol;"

"n(H_2)=0.866mol;"

"n(NH_3)=4.17\sdot10^{-8}mol;"

"V=4.00L;"

"N_2+3H_2\leftrightarrow 2NH_3;" "K=\dfrac{[NH_3]^2}{[N_2][H_2]^3};"

"[NH_3]=\dfrac{n(NH_3)}{V};"

"[NH_3]=\dfrac{4.17\sdot10^{-8}mol}{4.00L}=1.0425\sdot10^{-8}M;"

"[N_2]=\dfrac{n(N_2)}{V};"

"[N_2]=\dfrac{1.20mol}{4.00L}=0.3M;"

"[H_2]=\dfrac{n(H_2)}{V};"

"[H_2]=\dfrac{0.866mol}{4.00L}=0.2165M;"

"K=\dfrac{(1.0425\sdot10^{-8}M)^2}{0.3M\sdot(0.2165M)^3}=1.673\sdot10^{-15}";

Answer: "K=1.673\sdot10^{-15}."